OpenStudy (anonymous):

simplify the equations 1. 7xy - 2xz + 4xz - 12xyz + 2zx 2. -3(2x - 5) - 4(x + 7)

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

whoops I got it sorry. v.v

OpenStudy (anonymous):

haha lol

OpenStudy (anonymous):

1. 7xy-2xz+4xz-12xyz+2zx=7xy-12xyz+4xz=x(7y-12yz+4z) 2. -3(2x-5)-4(x+7)=-6x+15-4x-28=-10x-13=- (10x+13) like this?

OpenStudy (anonymous):

No combine all the xys and the xz first